OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth.

OpenStudy (hayleymeyer):

oh! i love this kind of stuff!!

OpenStudy (hayleymeyer):

Diameter = 8" Radius = 8"/2 = 4" Height = 3*4" = 12" Volume = πr 2 h=3.14⋅(4") 2 ⋅12"=3.14⋅192=602.88

OpenStudy (mrnood):

Yeah - well you may love it - but writing the whole solution is A) Against the code of conduct for this site B) Almost no help to the asker

OpenStudy (hayleymeyer):

oh, im sorry! :(

OpenStudy (anonymous):

Since Diameter = 2* Radius, it is 4". No more just "thinking". Substitute with "knowing". Okay, Diameter = 8" Radius = 8"/2 = 4" Height = ??

OpenStudy (anonymous):

It helps if you write stuff out. This allow you to check for errors and to explain and understand your work. Diameter = 8" Radius = 8"/2 = 4" Height = 3*4" = 12" Volume = πr2h=3.14⋅(4")2⋅12"=3.14⋅192=602.88

OpenStudy (anonymous):

? If this is the case, you should have gotten 603.18578948924, which, when rounding to the hundreths place is 603.19.